# R-PINN: Recovery-type a-posteriori estimator enhanced adaptive PINN

### v1

@@ 4.5 Wave equation
For this test, we consider a wave equation as follow,
$\begin{split}&\frac{\partial^{2}u}{\partial t^{2}}-3\frac{\partial^{2}u}{\partial x^{2}}=0,\quad(t,x)\in[0,6]\times[-5,5],\\ &u(0,x)=\frac{1}{\cosh(2x)}-\frac{0.5}{\cosh(2(x-10))}-\frac{0.5}{\cosh(2(x+10))},\\ &\frac{\partial u}{\partial t}(0,x)=0,\\ &u(t,-5)=u(t,5)=0,\end{split}$ (4.7)
@@
$u(t,x)=\frac{0.5}{\cosh(2(x-\sqrt{3}t))}-\frac{0.5}{\cosh(2(x-10+\sqrt{3}t))}+\frac{0.5}{\cosh(2(x+\sqrt{3}t))}-\frac{0.5}{\cosh(2(x+10-\sqrt{3}t))}.$ (4.8)
@@
standard PINN
(c) *
Fig. 16: The predicted solution (left), predicted error (middle) and the distribution of adaptive collocation points (right) of three methods solving the wave equation.
Fig. 17: The relative $L^{2}$ errors of three methods in the optimization process of solving the Poison equation the wave equation.
R-PINN FI-PINN standard PINN
$N_{\text{iter}}$ Relative $L^{2}$ $L^{\infty}$ $N_{2}$ Relative $L^{2}$ $L^{\infty}$ $L^{2}$ Relative $L^{2}$ $L^{\infty}$
$1$ $0.0572$ $0.0761$ $100$ $0.1110$ $0.1590$ $500$ $0.3448$ $0.4554$
$2$ $0.0076$ $0.0098$ $100$ $0.0317$ $0.0491$ $1000$ $0.3158$ $0.4214$
Table 4: The relative $L^{2}$ error and $L^{\infty}$ error by three methods solving the wave equation.
In this test, the R-PINN method quickly reaches its performance limit. Therefore, we adopt a single adaptive iteration process ($M=2$) and pre-train the network for 5000 iterations. We show the relative $L^{2}$ errors and $L^{\infty}$ errors in the Table 4. The relative $L^{2}$ error and $L^{\infty}$ error can reach $0.0076$ and $0.0098$ with R-PINN, which are obviously lower than the other two methods. The detailed decreasing process is convenient to be seen in Fig. 17 The exact solution is depicted in Fig. 15. The left panel of Fig. 16 compares the predicted solutions of the three methods. The middle panel of Fig. 16 illustrates the absolute errors across the domain, showing that the errors associated with R-PINN are smaller than those with FI-PINN and standard PINN. Finally, the right panel of Fig. 16 displays the adaptive collocation points for two adaptive methods and random points for the standard PINN. From the right panel, we can see that the same problem occurs for FI-PINN when dealing with multiple large errors region, so it can’t handle well.
@@ 5 Conclusion
In this paper, we present recovery-type a-posteriori estimator enhanced adaptive PINN that combines the PINN with a recovery type estimator which is popularly used in the adaptive finite element method. Our innovation lies in the use of the recovery-type a-posteriori error estimator, which is different from the residual of PINNs, and the design of a tailored sampling method R-AD. We test the performance of R-PINN in several PDE problems with sharp or singular solutions, which normal PINN struggles to handle effectively. The results demonstrate that R-PINN achieves significant improvements over normal PINN. More studies can be done to analyze the performance of R-PINN for more complex problems and we shall focus on and analyze these issues in our future work.

### v2

@@ 4.5 Wave equation
For this test, we consider a wave equation as follows,
$\begin{split}&\frac{\partial^{2}u}{\partial t^{2}}-3\frac{\partial^{2}u}{\partial x^{2}}=0,\quad(t,x)\in[0,6]\times[-5,5],\\ &u(0,x)=\frac{1}{\cosh(2x)}-\frac{0.5}{\cosh(2(x-10))}-\frac{0.5}{\cosh(2(x+10))},\\ &\frac{\partial u}{\partial t}(0,x)=0,\\ &u(t,-5)=u(t,5)=0,\end{split}$
@@
$u(t,x)=\frac{0.5}{\cosh(2(x-\sqrt{3}t))}-\frac{0.5}{\cosh(2(x-10+\sqrt{3}t))}+\frac{0.5}{\cosh(2(x+\sqrt{3}t))}-\frac{0.5}{\cosh(2(x+10-\sqrt{3}t))}.$
@@
Fig. 16: The predicted solutions (left) and the distribution of adaptively distributed points (right) of two adaptive sampling methods solving the wave equation.
Fig. 17: Comparison of absolute errors distributions across the domain for three different PINNs: R-PINN (left), FI-PINN (middle), and RBA-PINN (right).
Fig. 18: The relative $L^{2}$ error by three different PINNs in the optimization process of solving the wave equation.
In this test, we adopt the adaptive iteration process with the number of $M=4$ and pre-train the network for 5000 iterations. We show the relative $L^{2}$ error and $L^{\infty}$ error in the Table 4. The relative $L^{2}$ error and $L^{\infty}$ error can reach 0.0076 and 0.0098 with R-PINN, which are obviously lower than the other two methods. The decreasing trend of the relative $L^{2}$ error during training can be clearly observed in Fig. 18. The exact solution is depicted in Fig. 15. The left panel of Fig. 16 compares the predicted solutions of the two adaptive sampling methods. Fig. 17 illustrates the absolute errors across the domain, showing that the errors associated with R-PINN are smaller than those with FI-PINN and RBA-PINN. Finally, the right panel of Fig. 16 displays the adaptively distributed points for two adaptive sampling methods. From the right panel, it can be observed that FI-PINN encounters similar difficulties when resolving multiple regions with large errors, indicating its limited effectiveness in such scenarios, where R-PINN demonstrates better performance.
Fig. 19: Comparison of the relative $L^{2}$ error (left) and $L^{\infty}$ error (right) under varying mesh sizes for different values of $N_{2}$.
4.6 Tests on the effects of sampling and mesh
For R-PINN, two critical factors influencing accuracy are the number of collocation points and the mesh size. To evaluate their effects, we conduct experiments on the Poisson’s equation presented in Section 4.2 and examine how these two aspects affect the performance of R-PINN.
The collocation points consist of two components: the background collocation points and the adaptively distributed points. We begin by analyzing the combined influence of $N_{2}$ and the mesh size. Specifically, we fix $N_{1}=8000$ and vary the mesh size among $10\times 10$, $30\times 30$, $50\times 50$, and $70\times 70$, while also testing $N_{2}=25,\ 50,\ 100,\ 200$. The finest mesh used is $70\times 70$, which already generates 9800 elements. A finer mesh would result in only a small number of elements containing background collocation points, thus undermining the effectiveness of the method.
We first examine the effect of mesh size, illustrated in Fig. 3. A decreasing trend in the $L^{\infty}$ error is clearly observed with finer meshes, whereas the relative $L^{2}$ error does not follow a consistent trend. This is reasonable because the mesh is primarily used to guide a-posteriori error estimation and sampling, rather than to enrich the representational capacity of the solution space. Next, we analyze the impact of $N_{2}$, which can also be seen in Fig. 19 by comparing four different curves in each subfigure. Both the relative $L^{2}$ error and the $L^{\infty}$ error generally decrease as $N_{2}$ increases, with the exception of a significant increase in relative $L^{2}$ error when $N_{2}$ increases from 100 to 200. This anomaly results from an imbalance between the background collocation points and adaptively distributed points, where excessive adaptively distributed points destroy the intended sampling strategy.
Fig. 20: Comparison of the relative $L^{2}$ error (left) and $L^{\infty}$ error (right) under varying values of $N_{1}$.
Finally, we examine the effect of the number of background collocation points $N_{1}$. With the mesh size fixed at $50\times 50$ and $N_{2}=100$, we evaluate the error under different values of $N_{1}$, as shown in Fig. 20. It is clearly observed that both the relative $L^{2}$ and $L^{\infty}$ errors decrease with increasing $N_{1}$, indicating that more background points contribute to a more accurate solution.
@@ 5 Conclusion
In this paper, we present recovery-type a-posteriori estimator enhanced adaptive PINN that combines the PINN with a recovery type estimator which is popularly used in the adaptive finite element method. Our innovation lies in the use of the recovery-type a-posteriori error estimator, which is different from the residual of PINNs, and the design of a tailored sampling method RecAD. We compare the performance of the proposed R-PINN with FI-PINN and RBA-PINN on several problems featuring sharp gradients or singularities, which vanilla PINNs struggle to resolve effectively. The results demonstrate that R-PINN achieves significant improvements. More studies can be done to analyze the performance of R-PINN for more complex problems and we shall focus on and analyze these issues in our future work.
6 Acknowledgment
JJ was partially supported by China Natural National Science Foundation (No. 22341302), the Fundamental Research Funds for the Central Universities, JLU (93Z172023Z05), and the Key Laboratory of Symbolic Computation and Knowledge Engineering of Ministry of Education of China housed at Jilin University. YL acknowledged the support of NSF-DMS (No. 2208499). CZ was partially supported by the Strategic Priority Research Program of the Chinese Academy of Sciences (No. XDB0640000) and the National Natural Science Foundation of China (No. 12571445). The authors wish to thank the anonymous reviewers for their constructive comments and valuable suggestions, which significantly improved the quality of this manuscript.